40 g of solute are present in 140 g of solution. What is the weight fraction of solute in solution? We dissolved 20 g of sugar in 70 g of water. What% of sugar solution was formed? From 50 g of salt we prepared a solution with a weight fraction of 0.2. What is the total weight of the solution and what is the volume of water used? How much g of sodium hydroxide do we need to prepare 300 g of a 12% solution?


1
Expert's answer
2021-03-04T06:26:15-0500

1. The weight fraction of the solution equals:

weight fraction = mass of a component / mass of the mixture = 40 g / 140 g = 0.29

2. Percentage of sugar in the solution equals:

weight fraction = mass of sugar / (total weight of the solution) = 20 g / (20 g + 70 g) = 20 g / 90 g = 0.22

3. The total weight of the solution:

total weight = mass of solute / weight fraction = 50 g / 0.2 = 250 g

The volume of water:

volume of water = total weight - mass of solute = 250 g - 50 g = 200 g

4. The weight of sodium hydroxide equals:

mass of solute = total weight × weight fraction = 300 g × 0.12 % = 36 g